OpenStudy (milo123):

MEDAL A cylinder and a cone have the same diameter: 8 inches. The height of the cylinder is 3 inches. The height of the cone is 18 inches. Use π = 3.14. What is the relationship between the volume of this cylinder and this cone? Explain your answer by determining the volume of each and comparing them. Show all your work.

OpenStudy (milo123):

Volume of the cylinder: 3.14 x r^2 x 3 = 150.72 Volume of cone: 1/3 x 3.14 x 4^2 x 3 = 50.24

OpenStudy (milo123):

@Hero please help with tthe rest of this:)

OpenStudy (milo123):

@quickstudent please help with the rest of this problem:)

OpenStudy (3mar):

May I help?

OpenStudy (milo123):

yes!:)

OpenStudy (3mar):

Thank you.

OpenStudy (milo123):

np..i did the volume of both, but don;t really understand what there is to compare

OpenStudy (3mar):

I think this would help you.

OpenStudy (3mar):

We know that the volume of the cone is one third of the vo,ume of the cylinder having the same base area and height. You can use this fact to compare "this cone's volume" with "this cylinder's one"

OpenStudy (3mar):

@milo123

OpenStudy (milo123):

sorry .. i'm here my computer was being slow

OpenStudy (3mar):

No problem.

OpenStudy (milo123):

ok thats a good one.

OpenStudy (3mar):

Great!! so What did you get?

OpenStudy (3mar):

@milo123 reach something!?

OpenStudy (wolf1728):

Volume of a cylinder = PI * r^2 * height Volume of a cone = (PI * r^2 * height)/3 SO, with equal height and equal radius, a cylinder will have 3 times the volume of a cone.

OpenStudy (wolf1728):

Cylinder height = 3 inches Cone height = 18 inches Cone is 6 times higher than cylinder (radius is the same). Is that enough help?